survey of 31 randomly selected students finds that they save a mean of $82 per semester by

using a website. Assume the date comes from a normal distribution and the sample standard deviation is $18 per month.


Confidence Interval: What is the 99% confidence interval to estimate the population mean? (Round your

answers to two decimal places.)


____< u < _____


1
Expert's answer
2022-05-12T15:18:49-0400

The critical value for "\alpha = 0.01" and "df = n-1 = 30" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.749996."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}},\bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(82- 2.749996\times\dfrac{18}{\sqrt{31}},"

"82+ 2.749996\times\dfrac{18}{\sqrt{31}})"

"=(73.11, 90.89)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "73.11 < \mu < 90.89

," which indicates that we are 99% confident that the true population mean "\mu"  is contained by the interval "(73.11, 90.89)."



"73.11 < \mu < 90.89"